
Probability Test Practice
Authored by Amanda Shears
Mathematics
8th Grade
CCSS covered
Used 4+ times

AI Actions
Add similar questions
Adjust reading levels
Convert to real-world scenario
Translate activity
More...
Content View
Student View
9 questions
Show all answers
1.
MULTIPLE CHOICE QUESTION
3 mins • 1 pt
A fair 17-sided die is shown. How many outcomes are in the sample space when rolling the die two times?
2
17
34
289
Answer explanation
When rolling a die twice, each roll has 17 possible outcomes. Therefore, the total outcomes are 17 (first roll) x 17 (second roll) = 289.
2.
MULTIPLE CHOICE QUESTION
3 mins • 1 pt
What is the theoretical probability of choosing two red marbles in a row?
1/6
2/6
1/36
2/36
Answer explanation
To find the probability of choosing two red marbles in a row, multiply the probability of choosing a red marble first (1/6) by the probability of choosing a red marble second (1/6). Thus, 1/6 * 1/6 = 1/36.
3.
OPEN ENDED QUESTION
3 mins • 1 pt
A fair 17 sided die is shown. Determine the theoretical probability of rolling a number that is divisible by 3 two times in a row.
Type your answer as a decimal rounded to the nearest thousandth.
Evaluate responses using AI:
OFF
Answer explanation
The numbers divisible by 3 on a 17 sided die are 3, 6, 9, 12, and 15. Therefore on the first roll there is a 5/ 17 chance and on the second roll there is a 5/17 chance. The probability of rolling one of these numbers on both roll is (5/17)x(5/17) or 25/289. As a percent rounded to the nearest tenth that is 8.7%.
4.
MULTIPLE CHOICE QUESTION
3 mins • 1 pt
Each letter of the work MISSISSIPPI is written on a card and put into a bag. What is the theoretical probability of choosing an "I", putting it back and then drawing a "P"?
6.6%
1.7%
2.5%
6.7%
Answer explanation
The word 'MISSISSIPPI' has 11 letters: 4 I's and 2 P's. The probability of drawing an I is 4/11, and for P is 2/11. Thus, the combined probability is (4/11) * (2/11) = 2/121 which is equal to 0.0661157025, which is approximately 6.6%.
Tags
CCSS.HSS.CP.B.7
5.
OPEN ENDED QUESTION
3 mins • 1 pt
Determine the experimental probability of P(R, A). Type your answer as a percentage rounded to the nearest tenth.
Evaluate responses using AI:
OFF
Answer explanation
To find the experimental P(R, A), divide the number of times (R,A) occurred (7) by the total number trials (64). Multiply by 100 to convert to a percentage, then round to the nearest tenth for the final answer.
Tags
CCSS.7.SP.C.6
6.
MULTIPLE CHOICE QUESTION
3 mins • 1 pt
If Elizabeth conducted a second experiment and spun the spinner twice 700 times, she would expect the new results to ...
have exactly the same experimental probabilities as in the first experiment.
have experimental probabilities that are closer to the theoretical probabilities than the first experiment.
have outcomes that are exactly 10 times the outcomes of the first experiment.
have experimental probabilities that are exactly twice as much as in the first experiment.
Answer explanation
In a larger sample size of 1400 spins, the experimental probabilities are likely to converge closer to the theoretical probabilities due to the law of large numbers, making this the correct choice.
7.
MULTIPLE CHOICE QUESTION
3 mins • 1 pt
If you roll a regular six-sided die 240 times, how many times should you expect to roll a 5?
40
80
120
200
Answer explanation
When rolling a six-sided die, the probability of rolling a 5 is 1/6. Therefore, in 240 rolls, you expect to roll a 5 about 240 * (1/6) = 40 times. Thus, the correct answer is 40.
Tags
CCSS.7.SP.C.6
Access all questions and much more by creating a free account
Create resources
Host any resource
Get auto-graded reports

Continue with Google

Continue with Email

Continue with Microsoft
or continue with
%20(1).png)
Apple
Others
By signing up, you agree to our Terms of Service & Privacy Policy
Already have an account?